OpenStudy (anonymous):

X+Y+Z=44 > x-Y-z=2 > x-y=7

OpenStudy (anonymous):

X+Y+Z=44 x-Y-z=2 x-y=7

OpenStudy (anonymous):

Can we assume that you are only using 3 variables and that capital letters are the same as lowercase?

OpenStudy (anonymous):

x=7+y(eq 3) 2x=46(by adding eq 1+2) x=23(same as above) 23-7=y y=16 23+16+z=44 z=5

OpenStudy (anonymous):

Add equations (1) and (2) to get 2x = 46 Thus x = 23 Thus, from (3) y = 16 and from 1, z = 5

OpenStudy (anonymous):

Can you do word problems on this website also?

OpenStudy (anonymous):

post anything you want people will answer

OpenStudy (anonymous):

Thank you for the answer.